Considering the following two projects, if the WACC =10%, then I should: Project Year 0 1...

  1. Considering the following two projects, if the WACC =10%, then I should:

Project

Year

0

1

2

3

4

A

Cash flows

-$100

$50

$60

$70

$80

B

Cash flows

-$100

$50

$50

$50

$50

  1. Accept project A since its NPV = $102.27 is higher than project B’s NPV = $58.49
  2. Accept project B since its NPV = $102.27 is higher than project A’s NPV = $58.49
  3. Project A’s NPV = - $102.27. Thus, I shall reject project A and accept project B
  4. Project B’s NPV = - $58.49. Thus, I shall reject project B and accept project A
  5. We should reject both projects since they produce negative NPVs.
Answer #1

Accept project A since its NPV = $102.27 is higher than project B’s NPV = $58.49

NPV of A=-100+50/1.1+60/1.1^2+70/1.1^3+80/1.1^4=102.2744

NPV of B=-100+50/1.1+50/1.1^2+50/1.1^3+50/1.1^4=58.4933